Given the following 4 scenarios: The contract interest rate was 3.5% and the expected inflation rate was 1.5%. The contract interest rate was 5% and the expected inflation rate was 2%. The contract interest rate was 7.5% and the expected inflation rate was 4%. The contract interest rate was 9% and the expected inflation rate was 5%. and an ex post actual inflation rate of 4.75%, answer both of the following questions. a) Indicate which scenario was expected to be best for the borrower and explain why it is best for the borrower. b) Indicate which scenario would have been best for society and explain why it is best for society. 3) Answer each if the following questions: a) If nominal GDP in 1996 was $8,100.2 billion, and 1996 real GDP in 2009 chained dollars was $10,561.0 billion, the 1996 GDP deflator index value was: b) If real GDP grows from $10,561.0 billion in 1996 to $11,034.9 billion in 1997, the growth rate for real GDP was: c) If the CPI is 156.9 in 1996 and 238.8 in 2016, then between 1996 and 2016, prices have increased by: (note: 1982-84 = 100)
Answer #1

Part-1) Scenario 1st is the best for lender because in this scenario the difference between the contract interest rate is 2% which is the ideal for the lender

PART-2) a) GDP deflator index value was:

(8,100.2/10561.0) x 100 = 76.6992

b) Growth rate for real GDP:

= (11,034,900,000 - 10,561,000,000) / 10,561,000,000

= 0.044873 * 100

= 4.49%

c) The prices have increased by:

= {238.8/156.9) - 1)}

= 52.19885%